i really need help on this question i keep checking my answer but it doesn't add up. {2(3-x)=16(x+1)}
Answer by lucky83(12)

{2(3-x)=16(x+1)}
6-2x=16x+16
6=18x+16
-10=18x
x=-10/18
x=-5/9